A $10,000 face value 12% coupon corporate bond matures on March 15, 2028. You purchase the bond on July 16, 2015 (M1) at a quoted price of 99.375. Please, compute the bond’s YTM, its invoice price as of the settlement date (M4), and its dirty price on the purchase date (M1). State the yield as a percentage with 6 digits after the decimal point and the prices in dollars and cents. Show all calculator inputs. You must use your calculator’s bond function and disregard weekends/holidays
Answer #1
Pv Quoted Price =Clean Price $9,937.50 (10000*0.99375)
Nper Number of years to mature 13 (2028-2015)
Pmt Annual coupon amount=12%*10000 $1,200
Fv Amount to be received on maturity $10,000
RATE Yield to maturity 12.097763% (Uing RATE function of excel with Nper=13, Pmt=1200, Pv=-9937.50,Fv=10000)
Dirty Price=Clean Price+Interest accrued till purchase date
Number of days from last coupon payment to purchase date 122 4 months=31+30+31+30)
Number of days from purchase date to next coupon payment 243 (365-107)
Accrued interest =Annual Coupon *(122/(122+243) $401.10
Dirty Price =9937.50+401.10 $10,338.60
Clean Price on M4:
Rate Yield to maturity 12.097763%
Nper Number of years to maturity 10 (13-3)
Pmt Annual coupon amount=12%*10000 $1,200
Fv Amount to be received on maturity $10,000
PV Clean Price on M4: $9,944.98 (Uing PV function of excel with Rate=10.097763%,Nper=10, Pmt=-1200, Fv=-10000)
Interest accrued $401.10
Invoice Price on M4=9944.98+401.10= $10,346.08